OpenStudy (anonymous):

Solve each inequality. Then graph the solution set on a number line. Help me pls! M-8>-12

jimthompson5910 (jim_thompson5910):

how would you solve for M

OpenStudy (anonymous):

I dont have an idea

jimthompson5910 (jim_thompson5910):

we want to move that -8 over so what can we do to both sides to do that?

OpenStudy (anonymous):

Put the 8 on both sides?

jimthompson5910 (jim_thompson5910):

subtract 8 from both sides? or add 8 to both sides?

OpenStudy (anonymous):

Subtract?

jimthompson5910 (jim_thompson5910):

-8 minus 8 = -16 but notice how -8 + 8 = 0

jimthompson5910 (jim_thompson5910):

so this means you must add 8 to both sides

OpenStudy (anonymous):

Wait im trying to gigure it out

jimthompson5910 (jim_thompson5910):

sorry

jimthompson5910 (jim_thompson5910):

just saw that message

OpenStudy (anonymous):

Can you msg it again I'm little confused about it

jimthompson5910 (jim_thompson5910):

M-8>-12 M-8+8 > -12+8 ... Add 8 to both sides. M + 0 > -4 M > -4 hopefully all that makes sense

jimthompson5910 (jim_thompson5910):

again, the reason why we add 8 to both sides is because -8+8 = 0 and that will go away on the left side

OpenStudy (anonymous):

12+8?

jimthompson5910 (jim_thompson5910):

it's -12+8

OpenStudy (anonymous):

Oh yeah sorry didnt see negative sign

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

I get it thanks but how about the graph solution

jimthompson5910 (jim_thompson5910):

to graph M > -4 you step 1) draw out a number line step 2) plot an open circle at -4 (it's open because there isn't a line under the > sign) step 3) shade to the right of the open circle you shade to the right because of the greater than sign

OpenStudy (anonymous):

I dont get it can you draw?

jimthompson5910 (jim_thompson5910):

which step doesn't make sense

OpenStudy (anonymous):

Okay nevermind. 6r<-36 my answer is r<-30 am i right?

jimthompson5910 (jim_thompson5910):

6r means 6 times r

jimthompson5910 (jim_thompson5910):

how do you undo multiplication?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

Why r is times? I dont get it

jimthompson5910 (jim_thompson5910):

that's what 6r means, it means 6 times r

OpenStudy (anonymous):

So its 0

OpenStudy (anonymous):

Then -36^6

jimthompson5910 (jim_thompson5910):

what operation is the opposite of multiplication

OpenStudy (anonymous):

Divide so i get -6

OpenStudy (anonymous):

R>-6?

jimthompson5910 (jim_thompson5910):

why did you flip the sign

jimthompson5910 (jim_thompson5910):

the inequality sign

OpenStudy (anonymous):

Oh yeah didnt see it

OpenStudy (anonymous):

What if the case is like this -w denominator is 4 <_ cant insert it -7

jimthompson5910 (jim_thompson5910):

what undoes division?

OpenStudy (anonymous):

Multiplication?

jimthompson5910 (jim_thompson5910):

good, so you multiply both sides by -4

OpenStudy (anonymous):

Did you get my question?

jimthompson5910 (jim_thompson5910):

yes you have to multiply both sides by -4

OpenStudy (anonymous):

I got 16 and 28 what is next?

jimthompson5910 (jim_thompson5910):

-1/4 times -4 is not 16

OpenStudy (anonymous):

How to use that 1/4 inncalculator?

jimthompson5910 (jim_thompson5910):

what is 1/4 in decimal form

OpenStudy (anonymous):

.25

jimthompson5910 (jim_thompson5910):

multiply that by 4

OpenStudy (anonymous):

-16?

OpenStudy (anonymous):

Its 16

jimthompson5910 (jim_thompson5910):

0.25 times 4 = ??

OpenStudy (anonymous):

1 or -1

jimthompson5910 (jim_thompson5910):

it's 1

OpenStudy (anonymous):

Then

jimthompson5910 (jim_thompson5910):

this means -w/4 <= -7 turns into 1w >= 28

jimthompson5910 (jim_thompson5910):

so the final answer is \[\large w \ge 28\]

OpenStudy (anonymous):

Ohh you can cancel the one

OpenStudy (anonymous):

Why the inequality signed change?

jimthompson5910 (jim_thompson5910):

when you multiply both sides by a negative number, the inequality sign flips

OpenStudy (anonymous):

Just multiply?

jimthompson5910 (jim_thompson5910):

and divide

OpenStudy (anonymous):

Oh tnx

OpenStudy (anonymous):

I really suck at math :(

jimthompson5910 (jim_thompson5910):

that's ok, just keep at it and you'll get better

OpenStudy (anonymous):

4x-15<_21 what does it mean if the inquality sign has sign at the bottom?

jimthompson5910 (jim_thompson5910):

it means that there's an "or equal to" part to it

OpenStudy (anonymous):

So see ifnim right ill mltiply the 4 and the x and ill put the 4 on both sides

jimthompson5910 (jim_thompson5910):

first you add 15 to both sides (to undo the -15) then you divide both sides by 4 (to undo the 4 times x)

OpenStudy (anonymous):

Ill put 15 and 4 to 21? -.-

OpenStudy (anonymous):

Cancel the 15

jimthompson5910 (jim_thompson5910):

\[\large 4x - 15 \le 21\] \[\large 4x - 15+15 \le 21+15\] \[\large 4x + 0 \le 36\] \[\large 4x \le 36\] \[\large \frac{4x}{4} \le \frac{36}{4}\] \[\large \frac{1x}{1} \le \frac{36}{4}\] \[\large 1x \le \frac{36}{4}\] \[\large x \le \frac{36}{4}\] \[\large x \le 9\]

OpenStudy (anonymous):

Woah every setting change you really need to memorize it

OpenStudy (anonymous):

Thanks i learned a lot from you ijust have one question on this number

OpenStudy (anonymous):

-16>_5(2z-11) ill times the 5 to the 2 and 11 right do i still need to bring down the 5 after I times it?

jimthompson5910 (jim_thompson5910):

\[\large -16 \ge 5(2z-11)\] \[\large -16 \ge 5*2z-5*11\] \[\large -16 \ge 10z-55\] I'll let you finish

OpenStudy (anonymous):

Thats what i did yeah

OpenStudy (anonymous):

Z>29 am i right?

jimthompson5910 (jim_thompson5910):

no that's incorrect

OpenStudy (anonymous):

-61 or -144

jimthompson5910 (jim_thompson5910):

first you add 55 to both sides

OpenStudy (anonymous):

Why and then

jimthompson5910 (jim_thompson5910):

what do you get when you add 55 to both sides

OpenStudy (anonymous):

-7.1

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

Aw i really dnt know

jimthompson5910 (jim_thompson5910):

\[\large -16 \ge 5(2z-11)\] \[\large -16 \ge 5*2z-5*11\] \[\large -16 \ge 10z-55\] \[\large -16+55 \ge 10z\] \[\large 39 \ge 10z\] I'll let you finish

OpenStudy (anonymous):

Z>49

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

10z means 10 times z

OpenStudy (anonymous):

So z<39

jimthompson5910 (jim_thompson5910):

that's also incorrect

OpenStudy (anonymous):

39>z

jimthompson5910 (jim_thompson5910):

divide both sides by 10

OpenStudy (anonymous):

Z<3.9

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Finally what does this mean define a variable and write an inequalityfor each problem

OpenStudy (anonymous):

Example the quotientof three times anumber and 4 is at least -16

jimthompson5910 (jim_thompson5910):

tell me what you got so far